40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4056     Entry time: Wed Dec 15 12:46:18 2010
Author: Koji 
Type: Summary 
Category: IOO 
Subject: Finishing up the vac work 

What else?

v: Edit on Dec 15 10PM
v: Edit on Dec 16 10PM

JD:  We should check OSEMs for all optics *after* table leveling.  Some of them (esp. BS and ITMX) are currently close to their limits right now.

KA: Check green alignment.

Take photos of the tables.

Fix the leveling weights



Location    Optics            Action
--------------------------------------------------------------
@ITMX -     v POX             alignment
            v POP1/POP2       alignment
            v Table Leveling

@ITMY -     POY               mirror replacement (45deg->0deg) / alignment
            v SR2-TT          alignment
            v SRM Tower       alignment / EQ-stop release
            v SRM             alignment
            v SRM OSEM
            vvSRM OPLEV (X2)  install (VIS)/ alignment
            v ITMY OPLEV (X2)   install (VIS)/ alignment
            v OM1/OM2         install (DLC 45deg)/ alignment       
            v Table Leveling

@BSC -      v OM3             install (DLC 45deg/ alignment)
            v OM4(PZT)          neutralize, adjustment
            IPPOS steering    alignment
            v BS OPLEV        alignment
           
v PRM OPLEV(x2)     alignment
            Beam dumps
            Table Leveling

@IMC -      v REFL              mirror replacement (45deg->0deg)

@ETMX -     Al foil removal
            Table Leveling

@ETMY -     ETMY damping
            OSEM
            OPLEV
            Al foil removal
            Table Leveling

@OMC -      v OM5(PZT)        neutralize, adjustment

@ITM/ETM -  Mirror Wiping

ELOG V3.1.3-